how to find the values of a trig. ratio for any value of angle. Say 2 degree or 35 or 56 or 11 etc ?? without calculator


"If the value is not given"


may i get any list of some angles n their ratio values which r commonly used in Ques ???

Grade:12

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear mohit ,

it is not possible to calculate all the values manually without the use of calc or table. Actually you need not bother about the calculation if table or calculator is not given as it will always be provided in the question then. nevertheless in case of specially the calculation is asked you can only calculate those values for which trigonometric identities like sin2x = 2 sinxcosx or sin3x=......   can be applied.  and also a  commonly used value is sin37 = 3/5  , cos37 = 4/5 , H.C verma uses these values very often.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum